OpenStudy (anonymous):

to graph the function g(x)=-6log_9x, first start with the graph f(x)=log_9x Now graph the function y=g(x) and use your graph to find the domain and range of g, and any asymtotes that g has

OpenStudy (anonymous):

\[f(x)=\log_9(x)\] really might as well be \[f(x)=\log(x)\] with only minor differences lets graph it

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=log_9%28x%29+domain+.1..20 goes through the points \((1,0)\) and \((9, 1)\) also if we went out that far \((81,2)\)

OpenStudy (anonymous):

domain of the log is \((0, \infty)\) i.e. positive numbers

OpenStudy (anonymous):

what does i.e mean

OpenStudy (anonymous):

range is \(\mathbb{R}\) or \((-\infty, \infty)\) or all real numbers i.e. "id est" which means simply "that is"

OpenStudy (anonymous):

it is not a math term, just something people say

OpenStudy (anonymous):

oh hahaha gotcha

OpenStudy (anonymous):

so then the retricemptote? x or y= and to what

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

assomtote

OpenStudy (anonymous):

god damn

OpenStudy (anonymous):

no idea how to spell that word lol

OpenStudy (anonymous):

asymptote

OpenStudy (anonymous):

how the hell does that turn into recipremete

OpenStudy (anonymous):

or whatever word

OpenStudy (anonymous):

because if you try to write A*S*S here it turns out like this retrice

OpenStudy (anonymous):

dumbbbb

OpenStudy (anonymous):

keeps you from saying 'you are an retriceole'

OpenStudy (anonymous):

well good thing for that because what an offensive word

OpenStudy (anonymous):

in any case lets see what \[g(x)=-6\log_9(x)\] looks like

OpenStudy (anonymous):

the domain is still the same all positive numbers and the range is also the same all real numbers

OpenStudy (anonymous):

but instead of increasing this one is decreasing because of the \(-6\) out front the asymptote is the \(y\) axis aka \(x=0\)

OpenStudy (anonymous):

you're explaining it so well I just do not understand :((((

OpenStudy (anonymous):

But that is the correct answer so thank you haha

OpenStudy (anonymous):

the picture tells it all

OpenStudy (anonymous):

in any case the domain of the log is always positive numbers and the range is always all numbers

OpenStudy (anonymous):

and your welcome now it is my bed time good luck with the rest

OpenStudy (anonymous):

okay thank you so much again for all of youre help tonight I seriously appreciate it so much!!!!! Sleep tight